What is the minimum number of consecutive even numbers that you must multiply together to be sure that the ones digit of their product is a 0?
a) 4
b) 5
c) 3
d) 2
e) 6

Found 2 solutions by checkley77, user_dude2008:
Answer by checkley77(12844)

THE ANSWER IS D) 2
0*2=0 OR
8*10=80

Answer by user_dude2008(1862)

answer is b) 5

worst case: 2*4*6*8*10 = 3840